# A Single Motor Nano Aerial Vehicle with Novel Peer-to-Peer Communication and Sensing Mechanism

### v1

@@ III-B Horizontal and Vertical Movement Experiment
In the movement experiments, instead of remaining at a position, MP3 moves between way points. The setup is shown in Fig. 13. The reference trajectories and the Y coordinates versus time in the 6 horizontal movement experiments are shown in Fig. 14. The reference trajectories and the Z coordinates versus time in the 6 vertical movement experiments are shown in Fig. 15. The experiments clearly show that MP3 is capable of controlled motion in 3D using the on-board communication and sensing system.
Figure 14: Reference trajectories of MP3 in 6 horizontal movement experiments. (a) and (b) shows the trajectory and (c) shows the Y-T plot where the thick red line is the target trajectory.
Figure 15: Reference trajectories of MP3 in 6 vertical movement experiments. (a) and (b) shows the trajectory and (c) shows the Z-T plot where the thick red line is the target trajectory.
@@ III-C Peer-to-Peer Communication and Sensing Experiment
In this set of experiments, we demonstrate the capability of MP3 to do peer-to-peer communication and sensing. The setup is shown in Fig. 16. Unlike in previous experiments, MP3 #1 is actively transmitting its current sensed position to MP3 #2, acting like a beacon. In software, we made sure that MP3 #1 can only communicate with the first three beacons, and MP3 #2 can only communicate with the last two beacons and MP3 #1. MP3 #2 will record two sets of position estimates, one using both the information from the beacons and the relative measurements and position estimates of MP3 #1, and another using only the information from the beacons. The horizontal reference trajectories of all 7 flights are in Fig. 17.
@@
We are especially interested in the flight performance of MP3 #2 horizontally, which most directly shows the effect of the information provided by MP3 #1. The localization error’s distribution in X and Y direction with or without peer-to-peer information is shown in Fig. 18(a) and (b). While the peer-to-peer communication has little impact on the Y direction’s localization accuracy, it reduces the RMS error of localization in X direction by approximately 50%. This result is consistent with the uncertainty given by (9).
@@ V Conclusion
In this paper, we presented MP3, a minimalist, single propeller drone equipped with novel peer-to-peer communication and sensing mechanisms. We explained the hardware and software architecture of the drone, especially its communication and sensing system, and demonstrated its capability to localize itself by communicating with and sensing its relative bearing, distance, and elevation to neighboring MP3s. We also showed that MP3 is capable of flying stably alone as well as together with peers. Thanks to the single-motor design and the novel communication and sensing system, MP3 is about 50% lighter than the lightest drone that has relative position sensing capabilities of similar accuracy [35] known to the authors. Having these three key capabilities, MP3 is theoretically capable to execute a wide variety of swarm tasks like collaborative search or shape formation in a fully distributed manner.
@@ Acknowledgments
I would like to thank Marko Vejnovic for reviewing the program’s architecture and providing helpful feedback.
This work was supported by The National Science Foundation, NRI2.0 grants 2024692 and 2024615

### v2

@@ III-B Horizontal and Vertical Movement Experiment
In the movement experiments, instead of remaining at a position, MP3 moves between way points. The setup is shown in Fig. 13. MP3’s reference trajectory and its Y coordinate versus time in the 6 horizontal movement experiments are shown in Fig. 14. MP3’s reference trajectory and its Z coordinate versus time in the 6 vertical movement experiments are shown in Fig. 15. The experiments clearly show that MP3 is capable of controlled motion in 3D using the on-board communication and sensing system.
Figure 14: Reference trajectories of MP3 in 6 horizontal movement experiments. (a) and (b) show the trajectory, and (c) shows the Y-T plot where the thick red line is the target trajectory.
Figure 15: Reference trajectories of MP3 in 6 vertical movement experiments. (a) and (b) show the trajectory, and (c) shows the Z-T plot where the thick red line is the target trajectory.
@@ III-C Peer-to-Peer Communication and Sensing Experiment
In this set of experiments, we demonstrate the capability of MP3 to do peer-to-peer communication and sensing. The setup is shown in Fig. 16. Unlike in previous experiments, MP3 #1 is actively transmitting its current sensed position to MP3 #2, acting like a beacon. In software, we made sure that MP3 #1 can only communicate with the first three beacons, and MP3 #2 can only communicate with the last two beacons and MP3 #1. MP3 #2 will record two sets of position estimates, one using both the information from the beacons and the relative measurements and position estimates of MP3 #1, and another using only the information from the beacons. The horizontal reference trajectory of MP3s in all 7 flights are shown in Fig. 17.
@@
We are especially interested in the flight performance of MP3 #2 horizontally, which most directly shows the effect of the information provided by MP3 #1. The localization error’s distribution in X and Y direction with or without peer-to-peer information is shown in Fig. 18(a) and (b). While the peer-to-peer communication has little impact on the localization accuracy in Y direction, it reduces the RMS error of localization in X direction by approximately 50%. This result is consistent with the uncertainty given by (9).
@@ V Conclusion
In this paper, we presented MP3, a minimalist, single propeller drone equipped with novel peer-to-peer communication and sensing mechanisms. We explained the hardware and software architecture of the drone, especially its communication and sensing system, and demonstrated its capability to localize itself by communicating with and sensing its relative bearing, distance, and elevation to neighboring MP3s. We also showed that MP3 is capable of flying stably alone as well as together with peers. Thanks to the single-motor design and the novel communication and sensing system, MP3 is about 50% lighter than the lightest drone known to the authors that has relative position sensing capabilities of similar accuracy [35]. Having these three key capabilities, MP3 is theoretically capable to execute a wide variety of swarm tasks like collaborative search or shape formation in a fully distributed manner.
@@ Acknowledgments
Thanks to Marko Vejnovic for reviewing the program’s architecture and providing helpful feedback. This work was supported by The National Science Foundation, NRI2.0 grants 2024692 and 2024615.